Please help me solve for x
4/7x=20
show work please because I'm clueless. Thank you

Answer by rm29924(97)

4/7x=20
to get rid of the fraction multiply both sides by 7
7 is the same as 7/1
7(4/7x)=7(20)
7/1(4/7x)=7(20)
28/7x=140
4x=140
divide both sides by 4 to get the variable x by itself
x=35
check work:
4/7x=20
4/7(35)=20
140/7=20
20=20